OpenStudy (loinducoeur):

How would you answer a coordinate geometry question that involved finding the equation of a line? I want to see what your method would be like compared to my own. There's an example question down there, somewhere, too.

OpenStudy (anonymous):

do you have an example we could work with?

OpenStudy (loinducoeur):

I can find one, if you want. I'm just curious as to how other people would do it. Wait a bit, and I'll get one.

OpenStudy (anonymous):

cool

OpenStudy (loinducoeur):

The points A and B have coordinates (-2, 10) and (12,3) respectively. The line, L, is perpendicular to AB and intersects the y-axis at the point C (0,-1). The lines AB and L intersect at the point D. Write down the equation of L.

OpenStudy (loinducoeur):

If it helps, I can give you the equation of AB.

OpenStudy (accessdenied):

I would typically use point-slope form of a line because slope is relatively easy to compute (or given or related to another slope by being parallel / perpendicular), and you only need to know one other point on the line to complete the equation. \( y - y_1 = m(x-x_1) \) \( m = \dfrac{y_2 - y_1}{x_2 - x_1} \) So, in your example question, I would first find the line equation for AB. First I need slope, so I compute that using the slope formula. Then plug in the slope and one point (probably (-2,10) for smallest possible numbers) to get the equation. The line L, I would then use the slope of AB to find the perpendicular slope. We are already given the point C, so that lines up for point-slope form as well (or slope-intercept; they both lead to the same place). After filling in that equation, I would have answered "Write down the equation of L.", so I would be done!

OpenStudy (loinducoeur):

Thank-you! :)

OpenStudy (accessdenied):

No problem! Always love some coordinate geometry from time to time. :P